OpenStudy (anonymous):

A school choir has 80 students; 36 of them are in seventh grade. Also, 40 students in the choir are girls; 18 of them are in seventh grade. What is the probability that a student picked at random from the choir is either a girl or is in seventh grade? A. 27/40 B. 29/40 C. 13/40 D. 33/40

hero (hero):

 \[P(A) \space\text{or} \space P(B) = \frac{18}{80} + \frac{18}{80} + \frac{22}{80}\]